\begin{thebibliography}{000000000}

%\bibitem[BT76]{BT76} E. Bedford and B. A. Taylor, {\it The Dirichlet problem for a complex Monge-Amp\`ere operator,} Invent. math. {\bf37} (1976)  1--44.

%\bibitem[BT77]{BT77} E. Bedford\ and\ B. A. Taylor,  {\it The Dirichlet problem for an equation of complex Monge-Amp\`ere type}, in Partial differential equations and geometry (Proc. Conf., Park City, Utah, 1977), 39--50, Lecture Notes in Pure and Appl. Math., 48, Dekker, New York. 

\bibitem[BT82]{BT82} E. Bedford and B. A. Taylor, 
{\it A new capacity for plurisubharmonic functions,} Acta Math. {\bf149} (1982)  1--40.

%\bibitem[BT87]{BT87} E. Bedford\ and\ B. A. Taylor, {\it Fine topology, \v Silov boundary, and $(dd\sp c)\sp n$.} J. Funct. Anal. {\bf 72} (1987), no.~2, 225--251.

%\bibitem[Be19]{Be19} R. J. Berman, {\it From Monge-Amp\`ere equations to envelopes and geodesic rays in the zero temperature limit}, Math. Z. {\bf 291} (2019), no.~1-2, 365--394. 



%\bibitem[BJZ]{BJZ} S. Benelkourchi, B. Jennane\ and\ A. Zeriahi, {\it Polya's inequalities, global uniform integrability and the size of plurisubharmonic lemniscates,} Ark. Mat. {\bf 43} (2005), 85--112. 


%\bibitem[BD12]{BD12}{R. Berman and J.-P. Demailly, \it Regularity of plurisubharmonic upper envelopes in big cohomology classes. \rm Proceedings of the Symposium ``Perspectives in Analysis, Geometry and Topology'' in honor of Oleg Viro (Stockholm University, May 2008), Eds. I.~Itenberg, B.~J\"oricke, M.~Passare, Progress in Math.\ {\bf 296}, Birkh\"auser/Springer, Boston (2012) 39--66.}


%\bibitem[BEGZ]{BEGZ} S. Boucksom, P. Eyssidieux, V. Guedj, A. Zeriahi, {\it Monge-Amp\`ere equations in big cohomology classes.} Acta. Math. {\bf 205}, (2010), 199--262.

%\bibitem[Bl93]{Bl93} Z. B\l ocki, {\it Estimates for the complex Monge-Amp\`ere operator.} Bull. Polish Acad. Sci. Math. {\bf 41} (1993)   151--157.


%\bibitem[Bl09]{Bl09} Z. B\l ocki, {\it On geodesics in the space of K\"{a}hler metrics}, in  Advances in geometric analysis, 3--19, Adv. Lect. Math. (ALM), 21, Int. Press, Somerville.

%\bibitem[BK07]{BK07} Z. B\l ocki,  S. Ko\l odziej, {\it On regularization of plurisubharmonic functions on manifolds,} Proc. Amer. Math. Soc. {\bf 135} (2007) 2089–2093.

%\bibitem[Bo12]{Bo12} S. Boucksom, {\it Monge-Amp\`ere equations on complex manifolds with boundary}, in  Complex Monge-Amp\`ere equations and geodesics in the space of K\"{a}hler metrics, 257--282, Lecture Notes in Math., 2038, Springer, Heidelberg.

%\bibitem[CKNS85]{CKNS85}L. Caffarelli, J. Kohn,  L. Nirenberg, J. Spruck, {\it The Dirichlet problem for nonlinear second-order elliptic equations. II. Complex Monge-Amp\`ere, and uniformly elliptic, equations,} Comm. Pure Appl. Math. {\bf 38} (1985), 209--252.

%\bibitem[Ce84]{Ce84} U. Cegrell,  {\it On the Dirichlet problem for a complex Monge-Amp\`ere operator,} Math. Z. {\bf 185} (1984), 247 - 251.

\bibitem[Ce98]{Ce98} U. Cegrell,  {\it Pluricomplex energy,} Acta Math. {\bf 180:2} (1998) 187-217.

%\bibitem[Ce04]{Ce04} U. Cegrell, {\it The general definition of the complex Monge-Amp\`ere operator,} Ann. Inst. Fourier (Grenoble), {\bf 54} (2004)  159-179.

\bibitem[CK06]{CK06}
U. Cegrell\ and\ S. Ko\l odziej, {\it The equation of complex Monge-Amp\`ere type and stability of solutions}, Math. Ann. {\bf 334} (2006), no.~4, 713--729. 

%\bibitem[CKZ11]{CKZ11} U. Cegrell, S. Ko\l odziej\ and\ A. Zeriahi, {\it Maximal subextensions of plurisubharmonic functions}, Ann. Fac. Sci. Toulouse Math. (6) {\bf 20} (2011), Fascicule Sp\'{e}cial, 101--122. 

%\bibitem[Ch00]{Ch00} X. Chen, {\it The space of K\"{a}hler metrics}, J. Differential Geom. {\bf 56} (2000) 189--234.

%\bibitem[CDS15]{CDS2} X. X.  Chen, S.  Donaldson and S.  Sun
%{\it K\"ahler-Einstein metrics on Fano manifolds. I: Approximation of metrics with cone singularities.} J. Amer. Math. Soc. {\bf 28}  (2015), 183--197. 

\bibitem[Ch87]{Ch87} P. Cherrier, {\it \'Equations de Monge-Amp\`ere sur les vari\'et\'es Hermitiennes compactes.} Bull. Sci. Math. {\bf 111} (2) (1987)  343--385.

%\bibitem[Chi13]{Chiose13} I. Chiose, {\it The K\"ahler rank of compact complex manifolds.} J. Geom. Anal. {\bf 26} (2016), no.~1, 603--615.

% \bibitem[CY80]{CY80} S-Y. Cheng and S-T. Yau,  {\it On the existence of a complete K\"ahler metric on noncompact complex manifolds and the regularity of Fefferman's equation}. Comm. Pure Appl. Math., {\bf 33} (1980),  507--544.

%\bibitem[CZ19]{CZ19} J. Chu\ and\ B. Zhou, {\it Optimal regularity of plurisubharmonic envelopes on compact Hermitian manifolds}, Sci. China Math. {\bf 62} (2019), no.~2, 371--380.

%\bibitem[Cz10]{Cz10} R. Czy\.{z}, {\it On a Monge-Amp\`ere type equation in the Cegrell class $\cE_\chi$}, Ann. Polon. Math. {\bf 99} (2010), 89--97.

%\bibitem [De82]{De82}{J.-P. Demailly, \it Estimations $L^2$ pour l'op\'erateur
%$\overline\partial$ d'un fibr\'e vectoriel holomorphe semi-positif au-dessus
%d'une vari\'et\'e k\"ahl\'erienne compl\`ete. \rm Ann.\ Sci.\ \'Ecole Norm.\
%Sup.\ 4e S\'er. \bf 15 \rm (1982), 457-511.}

%\bibitem [De94]{De94}{J.-P. Demailly, \it Regularization of closed positive currents of type $(1,1)$ by the flow of a Chern connection. \rm  Aspects of Math. Vol. E26, Vieweg (1994), 105--126.}

%\bibitem[DP04]{DP04} J-P. Demailly and M. Paun, {\it Numerical characterization of the K\"ahler cone of a compact K\"ahler manifold.} Ann. of Math. (2) {\bf 159} (2004), no. 3, 1247-1274.

%\bibitem [De85]{De85} J.-P. Demailly, {\it Measures de Monge-Amp\`ere et caract\'erisation g\'eom\'etrique des vari\'et\'es alg\'ebraiques affines.}  Mem. Soc. Math. France (N.S.) {\bf 19}  (1985),  1-124.



%\bibitem[DDGHKZ]{DDGHKZ} J.-P. Demailly, S. Dinew, V. Guedj, P. Hiep, S. Ko\l odziej and  A. Zeriahi, {\it H\"older continuous solutions to Monge-Amp\`ere equations.} J. Eur. Math. Soc. (JEMS) {\bf 16} (2014)  619--647.

%\bibitem[DF18]{DiF} E. Di Nezza\ and\ C. Favre, Regularity of push-forward of Monge-Amp\`ere measures, Ann. Inst. Fourier (Grenoble) {\bf 68} (2018), no.~7, 2965--2979.

%\bibitem[DL15]{DL15}E. Di Nezza\ and\ C. H. Lu, {\it Generalized Monge-Amp\`ere capacities}, Int. Math. Res. Not. IMRN {\bf 2015} 7287--7322.


%\bibitem[DL17]{DL17}E. Di Nezza\ and\ C. H. Lu, {\it Complex Monge-Amp\`ere equations on quasi-projective varieties}, J. Reine Angew. Math. {\bf 727} (2017)  145--167.



%\bibitem[Di09]{dinew09} S. Dinew, {\it An inequality for mixed Monge-Amp\`ere measures.} Math. Zeit. {\bf 262} (2009), 1--15.

\bibitem[Di16]{Di16} S. Dinew, {\it Pluripotential theory on compact Hermitian manifolds,} Ann. Fac. Sci. Toulouse Math. (6) {\bf 25} (2016), no.~1, 91--139.  

%\bibitem[Di19]{Di19} S. Dinew, {\it Lectures on Pluripotential theory on compact Hermitian manifolds,} Lecture notes in Mathematics {\bf 2246} (2019), "Complex Non-K\"ahler geometry", Springer.

\bibitem[DH12]{DH12}
S. Dinew\ and\ P.-H. Hiep, {\it Convergence in capacity on compact K\"{a}hler manifolds,} Ann. Sc. Norm. Super. Pisa Cl. Sci. (5) {\bf 11} (2012), no.~4, 903--919.


\bibitem[DK12]{DK12}S. Dinew and S. Ko\l odziej, 
{\it  Pluripotential estimates on compact Hermitian manifolds.}   Adv. Lect. Math. (ALM), {\bf 21} (2012)  International Press, Boston. 


%\bibitem[DZ10]{dinew-zhang10} S. Dinew and  Z. Zhang, {\it On stability and continuity of bounded solutions of degenerate complex Monge-Amp\`ere equations over compact K\"ahler manifolds.} Adv. Math. {\bf 225} (2010), no. 1, 367--388.

%\bibitem[DNS10]{DNS} T.-C. Dinh, V.-A. Nguyen, N. Sibony, {\it Exponential estimates for plurisubharmonic functions and stochastic dynamics,} J. Differential Geom. {\bf 84} (2010) 465-488. 

%\bibitem[DN14]{DN14} T.-C. Dinh, V.-A. Nguyen, {\it Characterization of Monge-Amp\`ere measures with H\"older continuous potentials.} J. Funct. Anal. {\bf 266} (2014), no. 1, 67--84.


%\bibitem[Do99]{Do99} S. K. Donaldson, {\it Symmetric spaces, K\"{a}hler geometry and Hamiltonian dynamics}, in  Northern California Symplectic Geometry Seminar, 13--33, Amer. Math. Soc. Transl. Ser. 2, 196, Adv. Math. Sci., 45, Amer. Math. Soc., Providence, RI. 

%\bibitem[EGZ09]{EGZ09} P. Eyssidieux, V. Guedj and A. Zeriahi, {\it Singular K\"ahler-Einstein metrics.} J. Amer. Math. Soc. {\bf 22} (2009),  607--639.

%\bibitem[FTWZ16]{FTWZ} S. Fang, V. Tosatti, B. Weinkove, T. Zheng, {\it Inoue surfaces and the Chern-Ricci flow,} J. Funct. Anal. {\bf 271} (2016), no.~11, 3162--3185.

%\bibitem[FY08]{FY08} J.-X. Fu and S.-T. Yau,  {\it The theory of superstring with flux on non-K\"ahler manifolds and  the complex Monge-Amp\`ere equation.} J. Differential Geom. {\bf 98} (2008), 369--428.

%\bibitem[FWW10a]{fww1} J. Fu, Z. Wang\ and\ D. Wu, {\it Form-type Calabi-Yau equations.} Math. Res. Lett. {\bf 17} (2010),  887--903. 

%\bibitem[FWW10b]{fww2} J. Fu, Z. Wang\ and\ D. Wu, {\it Form-type Calabi-Yau equations on K\"ahler manifolds of nonnegative orthogonal bisectional curvature.} Preprint, arXiv: 1010.2022.

%\bibitem[GS80]{GS80} T. W. Gamelin\ and\ N. Sibony, {\it Subharmonicity for uniform algebras}, J. Functional Analysis {\bf 35} (1980), no.~1, 64--108. 


%\bibitem[Gau77]{gauduchon77} P. Gauduchon, {\it La th\'eor\`eme de l'excentricit\'e nulle.} C. R. Acad. Sci. Paris {\bf 285} (1977), 387--390.

%\bibitem[GLZ19]{GLZ19}V. Guedj, C. H. Lu\ and\ A. Zeriahi, {\it Plurisubharmonic envelopes and supersolutions}, J. Differential Geom. {\bf 113} (2019), no.~2, 273--313.


%\bibitem[Gil11]{gill11} M. Gill, {\it Convergence of the parabolic complex Monge-Amp\`ere equation on compact Hermitian manifolds.} Comm. Anal. Geom. {\bf 19} (2011),  277--303.

%\bibitem[Gil13]{gill13} M. Gill, {\it The Chern-Ricci flow on smooth minimal models of general type.} Preprint arXiv: 1307.0066v1.

%\bibitem[Gu98]{Gu98} B. Guan, {\it The Dirichlet problem for complex Monge-Amp\`ere equations and regularity of the pluri-complex Green function}, Comm. Anal. Geom. {\bf 6} (1998)   687--703.

%\bibitem[GL10]{GL10} B. Guan and Q. Li, {\it Complex Monge-Amp\`ere equations and totally real submanifolds.} Adv. Math. {\bf 225} (2010)   1185--1223.

%\bibitem[GL21a]{GL21a} V. Guedj, C.-H. Lu, {\em Quasi-plurisubharmonic envelopes 2: Bounds on Monge-Amp\`ere volumes,} Preprint, {\bf arXiv:} 2106.04272.

%\bibitem[GL21b]{GL21b} V. Guedj, C.-H. Lu,  {\it Quasi-plurisubharmonic envelopes 3: Solving Monge-Amp\`ere equations on hermitian manifolds}, Preprint, {\bf arXiv:} 2107.01938. 


%\bibitem[GT]{GT98} D. Gilbarg\ and\ N. S. Trudinger, {\it Elliptic partial differential equations of second order}, reprint of the 1998 edition, Classics in Mathematics, Springer-Verlag, Berlin, 2001
%\bibitem[GZ05]{GZ1} V.Guedj and A.Zeriahi, {\it Intrinsic capacities on compact K\"ahler manifolds} J. Geom. Anal. {\bf 15} (2005),%607-639.



%\bibitem[GKZ08]{GKZ08} V. Guedj, S. Ko\l odziej, A. Zeriahi, {\it H\"older continuous solutions to Monge-Amp\`ere equations,} Bull. Lond. Math. Soc. {\bf 40} (2008)  1070-1080.

\bibitem[GL21a]{GL21a} V. Guedj, C.-H. Lu, {\em
Quasi-plurisubharmonic envelopes 2: Bounds on Monge-Amp\`ere volumes,}
Preprint, {\bf arXiv:} 2106.04272. to appear in J. Algebraic Geometry.

\bibitem[GL21b]{GL21b} V. Guedj, C.-H. Lu, {\em
Quasi-plurisubharmonic envelopes 3: Solving Monge - Amp\`ere equations on hermitian manifolds,} {\bf arXiv:} 2107.01938.


%\bibitem[GZ07]{GZ07} V. Guedj and A. Zeriahi, {\it The weighted Monge-Amp\`ere energy of quasiplurisubharmonic functions,} J. Funct. Anal. {\bf 250} (2) (2007), 442-482.

%\bibitem[GZ17]{GZ-book} V. Guedj\ and\ A. Zeriahi, {\it Degenerate complex Monge-Amp\`ere equations}, EMS Tracts in Mathematics, 26, European Mathematical Society (EMS), Z\"{u}rich, 2017.

%\bibitem[HQ19]{HQ19} L. M. Hai\ and\ V. V. Quan, H\"{o}lder continuity for solutions of the complex Monge-Amp\`ere type equation, J. Math. Anal. Appl. {\bf 494} (2021), no.~1, Paper No. 124586, 14 pp. 

%\bibitem[HQ21]{HQ21} L. M. Hai\ and\ V. V. Quan, {\it Existence and H\"{o}lder continuity to solutions of the complex Monge--Amp\`ere-type equations with measures vanishing on pluripolar subsets,} Internat. J. Math. {\bf 32} (2021), no.~14, Paper No. 2150099, 19 pp.

\bibitem[H08]{Hiep08}
P.-H.  Hiep, {\it On the convergence in capacity on compact Kahler manifolds and its applications,} Proc. Amer. Math. Soc. {\bf 136} (2008), no.~6, 2007--2018. 

%\bibitem[Hi10]{hiep} H.-H. Pham,  {\it H\"older continuity of solutions to the Monge-Amp\`ere equations on compact K\"ahler manifolds.} Ann. Inst. Fourier (Grenoble) {\bf 60} (2010), no. 5, 1857--1869.


%\bibitem[JY93]{JY93} J. Jost\ and\ S.-T. Yau, {\em A nonlinear elliptic system for maps from Hermitian to Riemannian manifolds and rigidity theorems in Hermitian geometry,} Acta Math. {\bf 170} (1993)   221--254.

%\bibitem[Ki78]{kis} C. O. Kiselman, {\it The partial Legendre transformation for plurisubharmonic functions.} Invent. math. {\bf 49} (1978),  137--148.

%\bibitem[Ko95]{Ko95} S. Ko\l odziej, {\em The range of the complex Monge-Amp\`ere operator. II,} Indiana Univ. Math. J. {\bf 44} (1995)  765--782. 

%\bibitem[Ko96]{Ko96} S. Ko\l odziej, {\it Some sufficient conditions for solvability of the Dirichlet problem for the complex Monge-Amp\`ere operator}, Ann. Polon. Math. {\bf 65} (1996)  11-21.


%\bibitem[Ko98]{Ko98} S. Ko\l odziej, {\it The complex Monge-Amp\`ere equation,} Acta Math. {\bf 180} (1998)  69-117.

%\bibitem[Ko00]{Ko00} S. Ko\l odziej, {\it Weak solutions of equations of complex Monge-Amp\`ere type}, Ann. Polon. Math. {\bf 73} (2000), no.~1, 59--67. 

%\bibitem[Ko03]{Ko03} S. Ko\l odziej,  {\it The Monge-Amp\`ere equation on compact K\"ahler manifolds.} Indiana Univ. Math. J. {\bf 52} (2003)   667--686.

\bibitem[Ko05]{K05} S. Ko\l odziej, {\it The complex Monge-Amp\`ere equation and pluripotential theory.} Memoirs Amer. Math. Soc. {\bf 178} (2005)  pp. 64.

%\bibitem[Ko08]{Ko08} S. Ko\l odziej,  {\it H\"older continuity of solutions to the complex Monge-Amp\`ere equation with the right hand side in $L^p$. The case of compact K\"ahler manifolds.} Math. Ann. {\bf 342} (2008)  379--386.

\bibitem[KN15]{KN1} S. Ko\l odziej and N.-C. Nguyen, {\it Weak solutions to the complex Monge-Amp\`ere equation on Hermitian manifolds.} Analysis, complex geometry, and mathematical physics: in honor of Duong H. Phong, Contemporary Mathematics, vol. 644 (American Mathematical Society,
Providence, RI, 2015)  141-158.



%\bibitem[KN16b]{KN3}S. Ko\l odziej\ and\ N. C. Nguyen, {\it Weak solutions of complex Hessian equations on compact Hermitian manifolds}, Compos. Math. {\bf 152} (2016), no.~11, 2221--2248.

\bibitem[KN18]{KN4}S. Ko\l odziej\ and\ N.-C. Nguyen,  {\it H\"older continuous solutions of the Monge-Amp\`ere equation on compact Hermitian manifolds,}  Ann. Inst. Fourier (Grenoble) {\bf 68} (2018)  2951--2964. 


\bibitem[KN19a]{KN2} S. Ko\l odziej\ and\ N.-C. Nguyen, {\it Stability and regularity of solutions of the Monge-Amp\`ere equation on Hermitian manifolds,} Adv. Math. {\bf 346} (2019)  264--304. 


%\bibitem[KN19b]{KN5} S. Ko\l odziej\ and\ N.-C. Nguyen, {\it A remark on the continuous subsolution problem for the complex Monge-Amp\`ere equation,}  Acta Math Vietnam (2019). https://doi.org/10.1007/s40306-019-00347-0

%\bibitem[KN6]{KN6} S. Ko\l odziej, N.-C. Nguyen, {\it An inequality between complex Hessian measures of H\"older continuous $m-$subharmonic functions and capacity.} Geometric Analysis, In Honor of Gang Tian's 60th Birthday, Progress in Mathematics series by Birkhauser. Editors: Jingyi Chen, Peng Lu, Zhiqin Lu, and Zhou Zhang,  to appear.

\bibitem[KN21a]{KN7} 
S. Ko\l odziej\ and\ N.-C. Nguyen, {\em Continuous solutions to Monge-Amp\`ere equations on Hermitian manifolds for measures dominated by capacity,} Calc. Var. Partial Differential Equations {\bf 60} (2021)  Paper No. 93, 18 pp. 

\bibitem[KN23a]{KN21}
S. Ko\l odziej\ and\ N.-C. Nguyen,{\em  The Dirichlet problem for the Monge--Amp\`ere equation on Hermitian manifolds with boundary,} Calc. Var. Partial Differential Equations {\bf 62} (2023), no.~1, Paper No. 1.

\bibitem[KN23b]{KN22}  
S. Ko\l odziej and\ N.-C. Nguyen, {\em Weak solutions to Monge--Amp\`ere type equations on compact Hermitian manifold with boundary,} J. Geom. Anal. {\bf 33} (2023), no.~1, Paper No. 15. 

\bibitem[KT21]{KT21} 
S. Ko\l odziej\ and\ V. Tosatti, {\em Morse-type integrals on non-K\"{a}hler manifolds,} Pure Appl. Math. Q. {\bf 17} (2021), no.~3, 991--1004. 


\bibitem[LPT21]{LPT20} C. H. Lu, T.-T. Phung\ and\ T.-D. T\^{o}, {\em Stability and H\"{o}lder regularity of solutions to complex Monge-Amp\`ere equations on compact Hermitian manifolds,} Ann. Inst. Fourier (Grenoble) {\bf 71} (2021), no.~5, 2019--2045. 

\bibitem[Ni17]{Ni17} X. Nie, {\it Weak solution of the Chern-Ricci flow on compact complex surfaces,} Math. Res. Lett. {\bf 24} (2017)  1819--1844. 


%\bibitem[Ng14]{Ng14} N.-C. Nguyen, {\it Weak solutions to the complex Hessian equation,} Ph.D thesis, Jagiellonian University 2014. 

\bibitem[Ng16]{Ng16} N.-C. Nguyen, {\it The complex Monge-Amp\`ere type equation on compact Hermitian manifolds and Applications.} 
Adv. Math. {\bf 286} (2016)  240-285.

%\bibitem[Ng17]{Ng17} N.-C. Nguyen, {\it On the H\"older continuous subsolution problem for the complex Monge-Amp\`ere equation,} Calc. Var. Partial Differential Equations {\bf 57} (2018)  Art. 8, 15 pp.

%\bibitem[Ng18]{Ng18} N.-C. Nguyen, {\it On the H\"{o}lder continuous subsolution problem for the complex Monge-Amp\`ere equation, II,} Anal. PDE {\bf 13} (2020)  435--453.

%\bibitem[Nie13]{nie13} X. Nie, {\it Regularity of a complex Monge-Amp\`ere equation on Hermitian manifolds.} Comm. Anal. Geom. {\bf 22} (2014), no. 5, 833--856.

%\bibitem[PSS12]{pss12} D. H. Phong, J. Song and  J. Sturm, {\it Complex Monge-Amp\`ere equations.} Surveys in differential geometry. Vol. XVII, 327-410, Surv. Differ. Geom., {\bf 17}, Int. Press, Boston, MA, 2012.


%\bibitem[Po13]{popovici13} D. Popovici, {\it Aeppli Cohomology Classes Associated with Gauduchon Metrics on Compact Complex Manifolds.} Bull. Soc. Math. France {\bf 143} (2015),  763--800.


%\bibitem[Se92]{Se92} S. Semmes, {\it Complex Monge-Amp\`ere and symplectic manifolds}, Amer. J. Math. {\bf 114} (1992) 495--550. 

%\bibitem[SY97]{SY} R. Schoen\ and\ S. T. Yau, {\it Lectures on harmonic maps}, Conference Proceedings and Lecture Notes in Geometry and Topology, II, International Press, Cambridge, MA, 1997. 

%\bibitem[ST07]{ST1} J. Song, and G. Tian, {\it The K\"ahler-Ricci flow on surfaces of positive Kodaira dimension.}  Invent. math. {\bf 170}, (2007), 609--653.

%\bibitem[ST12]{ST3} J. Song\ and\ G. Tian, {\it Canonical measures and K\"ahler-Ricci flow,} J. Amer. Math. Soc. {\bf 25} (2012),  303--353.


%\bibitem[ST17]{ST2} J. Song and G. Tian, {\it The K\"ahler-Ricci flow through singularities.}  Invent. math. {\bf 207}, (2017), 519--595. 


%\bibitem[SzTo11]{SzTo11} G. Sz\'ekelyhidi\ and\ V. Tosatti, {\it Regularity of weak solutions of a complex Monge-Amp\`ere equation.} Anal. PDE {\bf 4} (2011),  369--378.

%\bibitem[SzTW]{SzTW} G. Sz\'ekelyhidi, \ V. Tosatti, \ B. Weinkove, {\it Gauduchon metrics with prescribed volume form,} arXiv:1503.04491. [math.DG].

\bibitem[Td18]{T18}
T.-D. T\^{o}, {Regularizing properties of complex Monge-Amp\`ere flows II: Hermitian manifolds,} Math. Ann. {\bf 372} (2018), no.~1-2, 699--741.

%\bibitem[TZh06]{tian-zhang06} G. Tian and  Z. Zhang, {\it On the K\"ahler-Ricci flow on projective manifolds of general type.} Chinese Ann. Math. Ser. B {\bf 27} (2006),  179--192.

%\bibitem[To09]{To1} V. Tosatti, {\it Limits of Calabi-Yau metrics when the K\"ahler class degenerates.} J. Eur. Math. Soc. (JEMS) {\bf 11} (2009),  755--776. 

%\bibitem[To10]{To2} V. Tosatti, {\it Adiabatic limits of Ricci-flat K\"ahler metrics, J. Differential Geom.} {\bf 84} (2010),  427--453.

\bibitem[To15]{To15} V. Tosatti, {\em Non-K\"{a}hler Calabi-Yau manifolds,} in  Analysis, complex geometry, and mathematical physics: in honor of Duong H. Phong, 261--277, Contemp. Math., {\bf 644}, Amer. Math. Soc., Providence, RI.

%\bibitem[To19]{To19} V. Tosatti, {\it Regularity of envelopes in K\"{a}hler classes}, Math. Res. Lett. {\bf 25} (2018), no.~1, 281--289.

%\bibitem[TW10a]{TW10a} V. Tosatti and B. Weinkove,  {\it Estimates for the complex Monge-Amp\`ere equation on Hermitian and balanced manifolds.} Asian J. Math. {\bf 14} (2010),  19--40.

\bibitem[TW10]{TW10b} V. Tosatti and B. Weinkove, {\it The complex Monge-Amp\`ere equation on compact Hermitian manifolds.} J. Amer. Math. Soc. {\bf 23} (2010)  1187--1195.

%\bibitem[TW12]{TW12}
%V. Tosatti\ and\ B. Weinkove, {\it Plurisubharmonic functions and nef classes on complex manifolds,} Proc. Amer. Math. Soc. {\bf 140} (2012), no.~11, 4003--4010.

%\bibitem[TW13]{TW12b} V. Tosatti and B. Weinkove,  {\it The Chern-Ricci flow on complex surfaces.} Compos. Math. {\bf 149} (2013), no. 12, 2101--2138.

%\bibitem[TW15]{TW12a} V. Tosatti and B. Weinkove, {\it On the evolution of a Hermitian metric by its Chern-Ricci form.} J. Differential Geom. {\bf 99} (2015),  125--163.

%\bibitem[TW17]{TW13} V. Tosatti and B. Weinkove,  {\it The Monge-Amp\`ere equation for $(n-1)$-plurisubharmonic functions on a compact K\"ahler manifold.} J. Amer. Math. Soc. {\bf 30} (2017),  311--346.

%\bibitem[TWY15]{TWYang15} V. Tosatti and B. Weinkove and W. Yang, {\it Collapsing of the Chern-Ricci flow on elliptic surfaces.}  Math. Ann. {\bf 362} (2015),  1223--1271.


%\bibitem[To18]{To18} D. T\^{o}, {\it Regularizing properties of complex Monge-Amp\`ere flows II: Hermitian manifolds,} Math. Ann. {\bf 372} (2018), no.~1-2, 699--741.

%\bibitem[Vu16]{viet16} D.-V, Vu {\it Complex Monge-Amp\`ere equation for measures supported on real submanifolds,} preprint, arXiv:1608.02794. to appear in Math. Ann

%\bibitem[Vu19]{viet19} D.-V. Vu, {\it Equilibrium measures of meromorphic self-maps on non-K\"{a}hler manifolds,} Trans. Amer. Math. Soc. {\bf 373} (2020), no.~3, 2229--2250.


%\bibitem[Yau76]{Y} S.-T. Yau, {On the Ricci curvature of a compact K\"ahler manifold and the complex Monge-Amp\`ere equation.} Comm. Pure Appl. Math. {\bf 31}   \rm (1978), 339--411.

%\bibitem[Zh17]{Zh17} T. Zheng, {\it The Chern-Ricci flow on Oeljeklaus-Toma manifolds,} Canad. J. Math. {\bf 69} (2017), no.~1, 220--240.

%\bibitem[Z21]{Z21} A. Zeriahi, {\it Remarks on the modulus of continuity of subharmonic functions}, \rm {\bf arXiv:} 2007.08399.

\end{thebibliography}

\documentclass{amsart}
\usepackage{amsmath, amsthm, amssymb}
\usepackage{verbatim}
\usepackage{tikz}
\usetikzlibrary{matrix}

\usepackage[mathscr]{eucal} 
\usepackage{mathrsfs} 
\usepackage{cancel}

\newcommand{\veps}{\varepsilon}
\newcommand{\vphi}{\varphi}

\newcommand{\al}{\alpha}

\newcommand{\la}{\lambda}
\newcommand{\om}{\omega}

\newcommand{\cP}{\mathcal{P}}
\newcommand{\cF}{\mathcal{F}}

\newtheorem{thm}{Theorem}
\newtheorem{prop}[thm]{Proposition}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{cor}[thm]{Corollary}

\theoremstyle{definition}

\newtheorem{remark}[thm]{Remark}

\numberwithin{thm}{section}
\numberwithin{equation}{section}
\renewcommand{\[}{\begin{equation}}
\renewcommand{\]}{\end{equation}}

\newcommand{\wed}{\wedge}

\title[Weak convergence of Monge-Amp\`ere measures]{Weak convergence of Monge-Amp\`ere measures on compact Hermitian manifolds}

\author{S\l awomir Ko\l odziej and Ngoc Cuong Nguyen} 
\address{Faculty of Mathematics and Computer Science, Jagiellonian University, \L ojasiewicza 6, 30-348 Krak\'ow, Poland}
\email{slawomir.kolodziej@im.uj.edu.pl}
\address{Department of Mathematical Sciences, KAIST, 291 Daehak-ro, Yuseong-gu, Daejeon 34141, South Korea}
\email{cuongnn@kaist.ac.kr}
\begin{document}

\maketitle

\begin{abstract}
We give a sufficient condition on a sequence of uniformly bounded $\om$-plurisubharmonic functions, $\om$ being a Hermitian metric, for which the sequence of associated Monge-Amp\`ere measures converges weakly. This criterion can be used to obtained a bounded $\om$-plurisubharmonic solution to the Monge-Amp\`ere equation.
\end{abstract}

\bigskip
\bigskip

\section{Introduction}

Let $(X, \om)$ be a compact Hermitian manifold of dimension $n$. The complex Monge-Amp\`ere equation  on those manifolds has been studied extensively in recent years. The classical solution in the smooth case was provided by Tosatti-Weinkowe \cite{TW10b} (for $n=2$ the equation was solved earlier by Cherrier \cite{Ch87}). The weak solutions to the equation were studied in \cite{DK12}, \cite{KN1, KN4, KN2, KN7}, \cite{LPT20} and recent advances for the semi-positive Hermitian form are contained in  \cite{GL21b}. The equation has found numerous geometric applications \cite{To15}, \cite{Ng16}, \cite{Di16}, \cite{Ni17}, \cite{T18} \cite{KT21} and \cite{GL21a}.

To obtain the weak solutions one often uses the stability of potentials of the approximating equations. 
It is well-known that the $L^p$-convergence of potentials does not imply the weak convergence of corresponding Monge-Amp\`ere measures. Furthermore, in the Hermitian setting there are fewer such criterions available as  compared to the K\"ahler manifolds.  In this note, we prove such a criterion  under suitable assumptions: of uniform boundedness and domination by Monge-Amp\`ere measures of another "nice" sequence. Recall  the Bedford-Taylor capacity: for a Borel subset $E\subset X$,
$$	cap_\om(E) := \sup\left\{\int_E (\om + dd^cw)^n : w\in PSH(X, \om), 0\leq w\leq 1\right\}.
$$
Here $PSH(X,\om)$ denotes the set of all $\om$-plurisubharmonic ($\om$-psh) functions on $X$. 
A sequence $\{\vphi_j\}_{j=1}^\infty \subset PSH(X,\om)$ is said converge in capacity to $\vphi\in PSH(X,\om)$ if for a given $\veps>0$, 
$$	
	\lim_{j\to +\infty} cap_\om(|\vphi_j -\vphi|>\veps) =0.
$$
The main result of this expository note is as follows.

\begin{thm} \label{thm:main} Let $\{u_j\}$ be a uniformly bounded sequence of $\om$-psh functions. Assume $(\om +dd^c u_j)^n \leq C (\om + dd^c \vphi_j)^n$  for some uniformly bounded sequence $\{\vphi_j\}$ such  that $\vphi_j \to \vphi \in PSH(X,\om)$ in capacity. If $u_j \to u \in PSH(X,\om) \cap L^\infty(X)$ in $L^1(X)$, then  a subsequence of $(\om+ dd^c u_j)^n$ converges weakly  to $(\om +dd^cu)^n$. 
\end{thm}
This is a generalization of \cite[Lemma~2.1]{CK06} from the local setting to compact Hermitian manifolds that is pointed out in the proof of  \cite[Lemma~2.11]{KN22} (see also \cite{KN21}). On compact K\"aher manifolds there are stronger results \cite[Theorem~2.1]{Hiep08} and \cite{DH12}. We will see that if the dimension $n=2$, then we have a similar result (Proposition~\ref{prop:cap-convergence}). However, we do not know if it still holds for dimensions $n\geq 3$.
It is also possible to extend the  theorem to  the case of Hermitian semi-positive $(1,1)$-forms $\al$ and we will consider it in a future  paper. 

An application of the theorem is that it provides a shorter proof of the existence of bounded $\om$-psh solutions to Monge-Amp\`ere equations with the right hand side in $L^p(X)$,  $p>1$.

\begin{cor} Let $0\leq f \in L^p(X)$ and $p>1$. Then, there exists a bounded function $u \in PSH(X,\om) \cap L^\infty(X)$ and a constant $c>0$ solving $\om_u^n = c f \om^n.$
\end{cor}

\begin{proof} Approximating $f_j \to f$ by smooth positive function $f_j$. By the Tosatti-Weinkove theorem \cite{TW10b} there are  $u_j \in PSH(X,\om) \cap C^\infty(X)$ and $c_j>0$ solving
$$
	(\om + dd^c u_j)^n = c_j f_j \om^n, \quad  \sup_X u_j =0.
$$
Using \cite[Eq. (5.12)]{KN1} we have $c_j \leq C_0$ with a uniform $C_0>0$. Therefore, by \cite{DK12} we have  $- C_1 \leq u_j \leq 0$.  By passing to a subsequence we may assume that $u_j \to u$ in $L^1(X)$ and $c_j \to c \geq 0$. Then, 
$$
	\int_{X} |u_j-u| (\om+ dd^c u_j)^n = \int_X |u-u_j| c_j f_j \om^n  \leq  C \|u_j -u\|_{L^1(X)}^\frac{1}{q},
$$
where $1/p+1/q =1$.
Hence, by Lemma~\ref{lem:weak-convergence} (below) there is a subsequence of $\{u_j\}$, for simplicity still denoted by $\{u_j\}$, such that $\om_{u_j}^n \to \om_u^n = c f \om^n$ weakly. Note that since $u$ is bounded,  it follows from \cite[Remark~5.7]{KN1} that $c>0$.
\end{proof}

Another application is proving the existence of bounded solutions for  a more general class of measures.
A positive Radon measure $\mu$ on $X$ is well dominated by capacity or belong to $\cF(X,h)$ (see \cite{K05})  if
\[
\label{eq:dominate}
	\mu(E) \leq F_h( cap_\om (E)),
\]
for any Borel set $E \subset X$, where $F_h(x) = x/h(x^{-\frac{1}{n}})$ for some increasing 
function $h : \mathbb R_+ \rightarrow (0, \infty ) $ satisfying
\[
\label{eq:admissible}
	\int_1^\infty \frac{1}{x [h(x) ]^{\frac{1}{n}} }  \, dx < +\infty.
\]

\begin{cor} Let $\mu$ be a positive Radon measure with $\mu(X) >0$. Assume that $\mu$ is well dominated by capacity. Then, there exists a bounded solution $u\in PSH(X,
\om)$ and $c>0$ solving 
$
	\om_u^n = c \mu.
$
\end{cor}

\begin{proof} It follows easily from the proof of \cite[Theorem~3.1]{KN21}. % that there exists an approximate sequence $\mu_\veps \to \mu$ weakly as $\veps \searrow 0$.
\end{proof}
\begin{remark} The solutions obtained in Corollary~2.7 and Corollary~2.8 are H\"older continuous and continuous, respectively. However, to get these statements, we need a stronger stability of solutions as in \cite{KN1, KN7}.
\end{remark}

\medskip

{\em Acknowledgement.} The first author is partially supported by  grant  no. \linebreak 2021/41/B/ST1/01632 from the National Science Center, Poland. The second author is  partially supported by the start-up grant G04190056 of KAIST and the National Research Foundation of Korea (NRF) grant  no. 2021R1F1A1048185.

\section{Proof of Theorem~\ref{thm:main} and convergence in capacity}

This section is devoted to  prove the main theorem.
Since all functions under consideration are uniformly bounded, by normalizing 
\[ \sup_X u_j = \sup_X \vphi_j =0,
\]
they belong to the  subset $\cP_0$ of $PSH(X,\om)$ given by
$$\cP_0 =\left\{  v\in PSH (X,\om) \cap L^\infty(X):  \sup_X v=0 \right\}.$$
Let us denote  $dV = \om^n$ the volume form of $X$.
We have the following result essentially due to Cegrell \cite{Ce98},  a detailed proof is  given in \cite[Lemma~2.1]{KN21}. 
\begin{lem} \label{lem:L1-norm-convergence}
Let $d\la$ be a finite positive Radon measure on $X$ vanishing on pluripolar sets. Suppose moreover that $\{u_j\} \subset  \cP_0$  converges $dV$-a.e. to  $u \in \cP_0$. Then there exists a subsequence  $\{u_{j_s}\} \subset \{u_j\}$ such that
$$  \lim_{j_s\to +\infty} \int_X u_{j_s} d\la = \int_X u d\la.$$
\end{lem}

\begin{comment}%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{proof} Since $d\la$ is a finite measure it follows that $\sup_{j} \int_{X} |u_j|^2 d\la < +\infty$. So there exists a subsequence $\{u_j\}$ weakly converging to $v \in L^2 (d\la)$. By the Banach-Saks theorem we can find a subsequence $u_{j_k}$ such  that 
$$	
	F_k= \frac{1}{k} (u_{j_1} + \cdots + u_{j_k}) \to v \quad\text{in } L^2(d\la)
$$
as $k \to +\infty$. Extracting a subsequence $\{F_{k_s}\}$ of $\{F_k\}$ we  get $F_{k_s} \to v$ a.e in $d\la$, and also  that $F_{k_s}$ converges a.e to $u$ with respect to the Lebesgue measure.  Therefore, $(\sup_{s>t} F_{k_s})^* \searrow u$ everywhere as $t\to +\infty$.
It follows that  there is a subsequence which we still denote by $\{u_j\}$ such that
$$
	\lim_{j\to \infty} \int_X u_j d\la = \int_X v d\la = \lim_{s\to \infty} \int_X F_{k_s} d\la = \lim_{t\to \infty} \int_X \sup_{s>t} F_{k_s}  d\la = \int_X u d\la.
$$
where the first identity used the decreasing convergence property; the second one used the $d\la$-a.e convergence, and  the last used the fact that $d\la$ does not charge  pluripolar sets. This completes the proof.
\end{proof}
\end{comment}%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

An immediate consequence is

\begin{cor}\label{cor:L1-convergence}  There exists a subsequence, still denoted by $\{u_j\}$, such that
$$ \lim_{j\to \infty} \int_X |u_j - u| d\la  =0.
$$
\end{cor}

\begin{proof}Applying Lemma~\ref{lem:L1-norm-convergence} twice to the sequences $\{u_j\}$ and $\max\{u_j, u\}$,  we have (still denoting by $\{u_j\}$  the resulting subsequence)
$$\lim_{j\to \infty} \int_X u_j d\la = \int_X u d\la, \quad \lim_{j\to \infty} \int_X \max\{u_j, u\} d\la = \int_X u d\la.$$
Since $\max\{u,u_j\} = (u_j+u + |u_j-u|)/2$, integrating both sides and using the previous equations we get that
 $u_j \to u$ in $L^1(d\la).$
\end{proof}

The next result  is a global analogue of \cite[Lemma~2.3]{KN22} which says that under the assumption of Theorem~\ref{thm:main}, the sequence converges uniformly in $L^1$-norm with respect to a family of measures.  On compact complex manifolds the Cegrell inequality will not be needed. For the reader convenience, we give details of the proof.

\begin{lem} \label{lem:uniform-L1-convergence} Let $\{w_j\}_{j=1}^\infty \subset \cP_0$ be a uniformly bounded sequence  that converges in capacity to $w\in \cP_0$. Assume   $\sup_j \int_X (\om+ dd^c w_j)^n \leq C_1$ for some $C_1>0.$ Then, 
$$
	\lim_{j\to \infty}  \int_X |u- u_j| (\om+ dd^c w_j)^k \wed \om^{n-k} = 0 \quad \text{for } k=0,1,...,n.
$$
\end{lem}

\begin{proof} The case $k=0$ is the assumption on $\{u_j\}$, and we only give here the proof of the last inductive step, the other steps are  the very similar. In the proof  that follows the constant $C>0$ is a  uniform constant depending only on $X, \om$ and the uniform bounds for $\|u_j\|_{L^\infty}$ and $\|w_j\|_{L^\infty}$, it may change from line to line.

Note that $|u-u_j| = (\max\{u, u_j\} - u_j) + (\max\{u, u_j\} - u)$. By quasi-continuity of $\om$-psh functions and Hartogs' lemma,  we have $\phi_j := \max\{u, u_j\} \to u$ in capacity. Fix $\veps>0$. Then,  when $j $ is large,
$$\begin{aligned}
	\int _{X} (\max\{u, u_j\} -u) (\om+dd^c w_j)^n 
&\leq \int_{\{|\phi_j -u|>\veps\}} (\om+dd^c w_j)^n + \veps \int_{X} (\om+dd^c w_j)^n \\
&\leq C cap_\om (|\phi_j -u|>\veps) + C_1 \veps.
\end{aligned}$$
Therefore, $\lim_{j\to \infty}  \int_X (\phi_j-u) (\om+dd^c w_j)^n =0$. Next, we have for $j>k$,
$$
\int_X (\phi_j -u_j) (\om+dd^c w_j)^n - \int_X (\phi_j - u_j) (\om+dd^c w_k)^n 
= \int_X (\phi_j - u_j) dd^c (w_j - w_k) \wed T ,
$$
where $T= T(j,k) = \sum_{s=1}^{n-1} \om_{w_j}^s \wed \om_{w_k}^{n-1-s}$. By integration by parts, 
$$\begin{aligned}
	\int_X (\phi_j - u_j) dd^c (w_j - w_k) \wed T 
&= \int_X (w_j - w_k) dd^c \left[(\phi_j -u_j) \wed T \right].
\end{aligned}
$$
Note that for $h= \phi_j -u_j$,
$$ \begin{aligned}
	dd^c (h T) 
&= 	dd^c h \wed T + dh \wed d^c T - d^c h \wed dT + h dd^c T  \\
&=	dd^c h \wed T + dh \wed d^c \om \wed T_1 - d^c h \wed d\om \wed T_2 \\
&=: S_0 + S_1 + S_2.
\end{aligned}$$
Here notice that $T_1$ and $T_2$ are positive currents, and similar type as of $T$.

We now estimate each term $S_0, S_1$ and $S_2$ separately. Firstly, for the term $S_0$, 
$$
	\int_X (w_j - w_k) dd^c (\phi_j-u_j) \wed T \leq \int_X |w_j -w_k| (\om_{\phi_j}+ \om_{u_j}) \wed T.
$$
Since $\|w_j \|_\infty, \|u_j\|_\infty \leq A$ in $X$, it follows that 
$$\begin{aligned}
	 \int_{X} |w_j - w_k| (\om_{\phi_j} + \om_{u_j}) \wed T 
&\leq  A \int_{\{|w_j -w_k| >  \veps\}}  (\om_{\phi_j} + \om_{u_j}) \wed T  \\ &\quad+ \veps\int_{\{|w_j -w_k| \leq \veps\}}  (\om_{\phi_j} + \om_{u_j}) \wed T  \\
&\leq  A^{n+1} cap_\om( |w_j -w_k| >  \veps)  + 	C \veps,
\end{aligned}$$
where the uniform bound for the second integral on the right hand side follows from uniform boundedness of the potentials (see. e.g., \cite[Proposition~2.3]{DK12}).
Since $w_j\to w$ in capacity, it follows that there exists $k_0$ such that for every $j>k\geq k_0$  the left hand side is less than $2C \veps$.

Secondly, for the term $S_1$, by the Cauchy-Schwarz inequality \cite[Proposition~1.4]{Ng16} in the Hermitian setting,
$$ \begin{aligned}
&	\left|\int_X (w_j - w_k) dh \wed d^c \om \wed T_1 \right|^2 \\
&\leq 	C \int_X |w_j-w_k| dh\wed d^c h \wed \om \wed T_1 \int_X |w_j-w_k| \om^2 \wed T_1 \\
&\leq 	C \int_X |w_j-w_k| \om^2 \wed T_1.
\end{aligned}$$
Therefore, by a similar argument as in the first case for the integral on the right hand side is bounded by $C\veps$ for every $j> k \geq k_0$ (we may increase $k_0$ if necessary).
 
Lastly, the term $S_2$ is estimated similarly as  $S_1$.
Thus, 
$$\begin{aligned}
\int_X (\phi_j -u_j) (\om+ dd^c w_j)^n 
&\leq  \int_X (\phi_j -u_j) (\om+ dd^c w_k)^n \\ 
&\quad +	\left|\int_X (\phi_j -u_j) \om_{w_j}^n - \int_X (\phi_j - u_j) \om_{w_k}^n \right| \\
&\leq  \int_X (\phi_j -u_j) (\om+ dd^c w_k)^n  + 2C \veps \\
&\leq \int_X |u-u_j| (\om+ dd^c w_{k})^n + 2 C \veps.
\end{aligned}$$
Fixing $k=k_0$ and applying Corollary~\ref{cor:L1-convergence} for $d\la = (\om+dd^c w_{k_0})^n$, we get that 
$$
	 \int_X (\phi_j -u_j) (dd^c w_j)^n  \leq (2C + 1) \veps \quad\text{for } j \geq k_1 \geq k_0.
$$
Since $\veps>0$ is arbitrary, the proof of the lemma is completed.
\end{proof}

The proof of Theorem~\ref{thm:main} is an immediate consequence of Lemma~\ref{lem:uniform-L1-convergence} and the following 
 weak convergence. 

\begin{lem} \label{lem:weak-convergence}
Suppose that 
\[\label{eq:energy-convergence}
	\lim_{j\to +\infty} \int_X |u_j -u| \om_{u_j}^n =0.
\]
Then, there exists a subsequence $\{u_{j_s}\}$ of $\{u_j\}$ such that $\om_{u_{j_s}}^n$ converges to $\om_u^n$ weakly. 
\end{lem}

\begin{proof} 
Let $A>0$ be such that  $-A \leq u_j, u \leq 0$.
By passing to a subsequence we may assume further that $u_j \to u$ a.e. in $X$ with respect to $dV$. 
Note that
 $$u = (\limsup_{j\to \infty} u_j)^* = \lim_{j\to \infty} (\sup_{\ell \geq j} u_\ell)^*.$$
Set 
\[\label{eq:hartogs-s}
	w_{j} =\max \{u_j, u-1/j\}.
\]
By the Hartogs lemma $w_j$ converges to $u$ in capacity. Therefore, by the convergence theorem in \cite{DK12} (see also \cite{BT82}),  
$
	\lim_{j\to \infty} \om_{w_j}^n = \om_u^n.
$
Thanks to Lemma~\ref{lem:uniform-L1-convergence} and the assumption $w_j \to u$ in capacity, we have
\[\label{eq:uniform-L1}
	\int_{X} |u_j -u| (\om + dd^c w_j)^n \to 0 \quad \text{as } j\to +\infty.
\]

Now we are ready to conclude the proof of the lemma. By \eqref{eq:energy-convergence} and \eqref{eq:uniform-L1} we can choose  a subsequence $\{u_{j_s}\} \subset \{u_j\}$ so that $$\int_X |u-u_{j_s}| (\om + dd^c u_{j_s})^n + \int_X |u-u_{j_s}|(\om + dd^c w_s)^n < 1/s^2.$$ 
Recall from \eqref{eq:hartogs-s} that $w_s = \max\{u_{j_s}, u-1/s\}$ and this implies
$$
	{\bf 1}_{\{u_{j_s} > u-1/s\}} (\om+ dd^c w_s)^n = {\bf 1}_{\{u_{j_s} > u-1/s\}} (\om + dd^c u_{j_s})^n.
$$ 
Therefore,  for $\eta \in C^\infty(X)$,
$$\begin{aligned}
	\left| \int_X \eta \om_{u}^n - \int_X \eta \om_{u_{j_s}}^n\right| 
&\leq  \left| \int_X \eta \om_{u}^n - \int_X \eta \om_{w_s}^n\right| + \left| \int_X \eta \om_{w_s}^n - \int_X \eta \om_{u_{j_s}}^n\right| \\
&\leq \left| \int_X \eta \om_{u}^n - \int_X \eta \om_{w_s}^n\right| + \left| \int_{\{u_{j_s} \leq u-1/s \}} \eta \om_{w_s}^n - \eta \om_{u_{j_s}}^n\right|.
\end{aligned}$$
The first term on the right hand side goes to zero as $\om_{w_s}^n \to \om_u^n$. It remains to estimate the second term. Firstly, by the choice of $\{u_{j_s}\} $ at the beginning of this proof,
$$\begin{aligned}
	\left|\int_{\{u_{j_s} \leq u-1/s \}} \eta \om_{u_{j_s}}^n \right| 
&\leq \|\eta\|_{L^\infty} \int_{\{u_{j_s} \leq u-1/s \}} \om_{u_{j_s}}^n \\
&\leq s \|\eta\|_{L^\infty} \int_X |u- u_{j_s}| \om_{u_{j_s}}^n  \leq \frac{1}{s}\|\eta\|_{L^\infty} \to 0 \quad\text{as } s\to +\infty.
\end{aligned}$$
Similarly, 
$$ \begin{aligned}
 \left|\int_{\{u_{j_s} \leq u-1/s\}} \eta \om_{w_s}^n \right| 
 &\leq \|\eta\|_{L^\infty} \left| \int_{\{u_{j_s}\leq  u-1/s\}} \eta \om_{w_s}^n\right|  \\
 &\leq s \|\eta\|_{L^\infty} \int_X |u- u_{j_s}| \om_{w_s}^n  \to 0 \quad\text{as } s\to +\infty.
\end{aligned}
$$
The last two estimates complete the proof of the lemma.  This also completed the proof of Theorem~\ref{thm:main}.
\end{proof}

We shall have a similar criterion for convergence in capacity \cite[Proposition~2.8]{KN21}. Let us recall \cite[Theorem~3.5]{DK12} and \cite[Lemmas~2.1, 2.2]{KN1}. 
Let $B>0$ be a uniform constant satisfying 
$$\begin{aligned}
&	-B\om^2 \leq 2n dd^c \om \leq B \om^2, \\
&	-B \om^3 \leq 4n^2  d\om \wed d^c \om \leq B\om^3.
\end{aligned}
$$
Then, for $\vphi, \psi \in PSH(X, \om) \cap L^\infty(X)$ satisfying $\sup_{\{\vphi<\psi\}} (\psi -\vphi) \leq 1$, 
\[ \label{eq:comparison-V1} \begin{aligned}
	\int_{\{\vphi<\psi\}} \om_\psi^n  &\leq \int_{\{\vphi < \psi\}} \om_{\vphi}^n 
	+ C_n \max\{1,B\}^n \sum_{k=0}^{n} \int_{\{\psi< \vphi\}} \om_{\vphi}^k \wed \om^{n-k},
\end{aligned}
\] 
where $C_n$ is a dimensional constant. If  the dimension $n=2$, then we have a better estimate, namely,
\[ \label{eq:comparison-V2} \begin{aligned}
	\int_{\{\vphi<\psi\}} \om_\psi^2  &\leq \int_{\{\vphi < \psi\}} \om_{\vphi}^2 
	+ C_n \max\{1,B\} \int_{\{\psi< \vphi\}} \om^2,
\end{aligned}
\]

\begin{prop} \label{prop:cap-convergence} Let $\{u_j\}$ be the sequence in Theorem~\ref{thm:main}. Then, $u_j$ converges to $u$ in capacity if and only if 
\[\label{eq:cap-convergence}
	\lim_{j\to +\infty} \int_X |u_j-u| \om_{u_j}^k \wed \om^{n-k} = 0, \quad k=0,...,n.
\]
In particular, if $n=2$, then the sequence converges in capacity.
\end{prop}

\begin{proof} Suppose that $u_j \to u$ in capacity. It follows from Lemma~\ref{lem:uniform-L1-convergence} that \eqref{eq:cap-convergence} holds for every $k=0, 1, ...,n$. Conversely, suppose that \eqref{eq:cap-convergence} holds true. Let  $\veps>0$ be fixed and without loss of generality we may assume that 
$
	-1 \leq u_j, u \leq 0.
$
We have
$$
	cap_\om(|u_j-u|>\veps) \leq cap_\om (u_j-u>\veps) + cap_\om (u-u_j >\veps).
$$
By Hartogs' lemma, $\max\{u_j, u\} \to u$ in capacity. Hence, $$\lim_{j\to +\infty} cap_\om (u_j-u >\veps) \leq \lim_{j\to +\infty} cap_\om(\max\{u_j,u\} -u >0) =0.$$
It remains to prove that $cap_\om(u-u_j>\veps) \to 0$ as $j\to +\infty$. Let $-1\leq \rho \leq 0$ be a function in $PSH(X,\om)$. Then,
$$
	\{u_j < u -\veps\} \subset \left\{u_j < (1-\frac{\veps}{4}) u + \frac{\veps}{4} \rho - \frac{3\veps}{4} \right\} \subset  \left \{ u_j < u - \frac{\veps}{2}\right\}
$$
Applying  \eqref{eq:comparison-V1} for $\vphi = u_j $ and $\psi = (1-\veps/4) u + \veps \rho/4 - 3\veps/4$ and then using the previous inclusions for domains of integration on the left hand side and on the right hand side, we have
$$\begin{aligned}
&	\int_{\{u_j < u-\veps\}} \left[\om 
+ (1-\frac{\veps}{4}) dd^c u + \frac{\veps}{4} dd^c  \rho \right]^n  \\
&\leq \int_{\{u_j< u - \veps/2\}} \om_{u_j}^n 
	 + C_n \max\{1,B\}^n \sum_{k=0}^{n} \int_{\{u_j< u-\veps/2\}} \om_{u_j}^k \wed \om^{n-k},
\end{aligned}$$
Since the integrand on the left hand side is larger than $(\veps/4)^n\om_\rho^n$, it follows that 
\[ \label{eq:cap-inequality} \begin{aligned}
	\left(\frac{\veps}{4} \right)^n cap_\om (u_j< u-\veps) &\leq \int_{\{u_j < u -\veps/2\}} \om_{u_j}^n \\ 
&\quad	+ C_n \max\{1,B\}^n \sum_{k=0}^{n} \int_{\{u_j< u-\veps/2\}} \om_{u_j}^k \wed \om^{n-k}.
\end{aligned}
\] 
Notice that for $k=0,...,n$ we have
$$
	\int_{\{u_j < u -\veps/2\}} \om_{u_j}^k \wed \om^{n-k}  \leq \frac{2}{\veps}  \int_X |u_j -u| \om_{u_j}^k \wed \om^{n-k} .
$$
Combining this inequality with \eqref{eq:cap-convergence} and \eqref{eq:cap-inequality} we obtain
$\lim_{j\to +\infty} cap_\om(u_j < u -\veps) =0$. The proof is completed.

Finally let us  finish the proof in the case $n=2$. Indeed, by \eqref{eq:comparison-V2} the  inequality corresponding to \eqref{eq:cap-inequality} reads
$$
	\left(\frac{\veps}{4} \right)^2 cap_\om(\{u_j < u-\veps\}) \leq \int_{\{u_j< u-\veps/2\}}\om_{u_j}^2 + C_n \max\{1,B\} \int_{\{u_j < u-\veps/2\}} \om^2.
$$
The right hand side is bounded by 
$$
	\frac{2}{\veps}\int_{X} |u_j-u| \om_{u_j}^2 + \frac{2C_n}{\veps} \max\{1,B\} \int_X |u_j- u| \om^2,
$$
which tends to zero as $j$ goes to infinity.
Since $\veps>0$ is fixed we conclude the convergence in capacity in dimension 2.
\end{proof}

\end{document}